PREX/CREX May 14, 2019 10:00am

From PREX Wiki
Jump to: navigation, search

Back to Main Page >> PREX/CREX_Teleconferences

previous meeting << >> following meeting

Logistic information

 BlueJeans calling instructions:
 Toll-Free Number (U.S.&  Canada):  888-240-2560
 International toll number:         408-740-7256
 Bluejeans CODE:                  628 981 532
 Bluejeans link: https://bluejeans.com/628981532
 Offline room:   CEBAF Center F324-325

Agenda

  1. Installation progress: Ciprian docDB363
    • CG: work proceeding. We are working with survey and alignment to make sure we have a little interferences as possible.
    • CG: target will be moved us by additional 5 cm to avoid interferences with the vacuum enclosures.
  2. Compton polarimeter
    • DG: struggling with the birefringence of the cavity. The cube that we used to measure the polarization in the cavity introduces some polarization effect. Will install amplifier in the afternoon
    • JCC: Spare base for PREX and making preparations for CREX
  3. Moller polarimeter
    • BH: New target survey will be done. New copper foil; installed.
  4. Beam Modulation
    • VO: repaired some coil connections and re-measured. Still having problems with hardware. More progress has been made on the instant reply analysis. PK: will followup later today
  5. BPM studies
    • CP: reproduced real beam problem with generator. Working with lab folks for solution that is least invasive.
  6. Integrating DAQ updates
    • PK: changed the CH crate last week. It needs to be recabled. We will re-check the VQWK modules in the crate. Next is changing the timing for the VQWK boards. We no longer have ADCX18 in the CH. We should think about how to prevent the temp excursions from happening. Other crates could also be affected.
  7. Parity analyzer:
    • PK: the japan regression is consistent with post-pan.
    • PK: we should gauge if we need mini-runs.
  8. LQ1 viton seals haplog3749
    • CG: Adam finished his analysis and the plan is to install a shield only on the left.
  9. Injector studies:
    • CP: next week we'll have 3x 8h shifts up to FC1. Will let everyone know. Rotation of cathode, pickup tests, proper voltages and chopper scan, set IAs for the other cells.
  10. Detectors:
    • DM: both arms frames for the VDCs have been installed. R main det assembly installed. Tomorrow Jack will install the GEM frames.
    • ChG: working without GEMs. L is working with MPDs; working on R. hope to finish by this weekend.
  11. Cross talk work:
    • CP: we are working on it and have a working group setup. The batteries upstairs don't see any pickup. The batteries in the hut see ppm level pickup.
    • KK: we should come up with a setup for a large pickup upstairs to see how it propagates to see how sensitive we are at the CH and HRSes.
  12. Discussion for next weeks:
    • RR: will present comparison next week

Present

In F324

Paul K, Caryn P, Devi A, Tao Y, Adam Z, Ciprian G,

Online

Alexandre C, Bob M, Cameron C, Chandan C, Dave G, David A, Dustin M, Eric K, Jinlong Z, Juliette M, KK, Paul S, Ryan R, Seamus R, Silviu D, Rich H, Amali P, Victoria O, Nilanga L, Juan Carlos C, Bill H

Cannot make it

Kent P